OpenStudy (anonymous):

Solve the recurrence relation: \(a_n = 6a_{n-1} - 8a_{n-2}; a_0 = 1, a_1 = 0 \)

OpenStudy (anonymous):

\(a_2 = -8, a_3 = -48, a_4 = -224\\ x^n = 6x^{n-1} - 8x^{n-2} \equiv x^n - 6x^{n-1} + 8x^{n-2} = 0\\ x^2 - 6x + 8 = 0; x = 6,4\) So now I should just be able to find constants to fulfill \(a_n = A6^n + B4^n\) ?

OpenStudy (anonymous):

Not sure how you got that second equaation

OpenStudy (anonymous):

is this linear algebra?

OpenStudy (xapproachesinfinity):

Me neither!

OpenStudy (anonymous):

Discrete math. For linear homogeneous recurrence relations with constant coefficients you're supposed to be able to find a solution of the form \(S_n = t^n\) which comes out to that.

OpenStudy (anonymous):

Yeah, just solve for \(A\) and \(B\) and see if it works

OpenStudy (anonymous):

ok yeah it works